3. You owe your credit company $50,000 and plan to make no payments on your credit card for 1.0 year. Your credit card company currently charges you an APR of 20.0%. If you make efforts to improve your credit score by 100 points, your credit card company will adjust your APR by 4.0%. As the result of you efforts, how much less interest would you pay over 1.0 year?

  1. $1,386

  2. $2,000

  3. $2,356

  4. $2,443

  5. $3,412
Answer #1

Answer:

Correct answer is:

$2356

Explanation:

Current charges an APR = 20.0%

Credit card interest charges are compounded monthly

Hence effective interest rate = (1 + APR / Number of compounding periods) Number of compounding periods - 1

= (1 + 20%/12) 12 - 1

= 21.9391%

The Interest payment would have been = 50,000 * 21.9391% = $10,969.55

Post adjustment APR = 20% - 4% = 16%

Effective interest rate = (1 + 16%/12) 12 - 1

= 17.2271%

The Interest payment = $50,000 * 17.2271% = $8,613.55

Reduction in interest over one year = $10,969.55 - $8,613.55 = $2,356

As such option C is correct and other options A, B, D and E are incorrect.